how much would an ad cost if it measured 8cm by 10cm?
the original measurements are 5cm by 12 cm costing $1500.00

Answer by checkley77(12844)

5*12=60 CM^2
1500/60=$25 PER CM^2
8*10=80 CM^2
80*25=$2,000 COST OF THE LARGER AD.